# THOR: Tool-Integrated Hierarchical Optimization via RL for Mathematical Reasoning

### v1

@@ 2.4 Self-Correction During Inference
During inference, our model follows the think-act-observe loop. To exploit immediate feedback from tool execution, we introduce a self-correction mechanism that dynamically rectifies erroneous reasoning steps, as shown in Figure 3(c). Specifically, when an action $a^{t}$ fails to execute, it indicates that both the action $a^{t}$ and its associated reasoning step $r^{t}$ are likely incorrect. To explore different solving paths, the model backtracks to $r^{t}$ and partitions it into a prefix $r^{t}_{\text{pre}}$ and a suffix $r^{t}_{\text{suf}}$, as previously described. Conditioned on the history up to $r^{t}_{\text{pre}}$, the model then regenerates a new reasoning suffix $\hat{r}^{t}_{\text{suf}}$ and a revised action $\hat{a}^{t}$, which achieves online error correction during inference.
The correction procedure can be repeated for up to $N_{\text{corr}}$ attempts. Importantly, each attempt only requires regenerating the suffix $\hat{r}^{t}_{\text{suf}}$ and the corresponding action $\hat{a}^{t}$, rather than the entire trajectory. Thus, the additional computational cost is minimal compared to the total cost.
@@ 3 Experiments
Table 1: Comparison with state-of-the-art methods on mathematical benchmarks, the best results are in bold and the second-best are underlined. †  denotes our reimplementation results of Avg@4. ‡  indicates results from their official releases.

### v2

@@ 2.4 Self-Correction During Inference
During inference, our model follows the think-act-observe loop. To exploit immediate feedback from tool execution, we introduce a self-correction mechanism that dynamically corrects erroneous reasoning steps, as shown in Figure 3(c). Specifically, when an action $a^{t}$ fails to execute, it indicates that both the action $a^{t}$ and its associated reasoning step $r^{t}$ are likely incorrect. To explore alternative solving paths, the model backtracks to $r^{t}$ and partitions it into a prefix $r^{t}_{\text{pre}}$ and a suffix $r^{t}_{\text{suf}}$, as previously described. Conditioned on the history up to $r^{t}_{\text{pre}}$, the model then regenerates a new reasoning suffix $\hat{r}^{t}_{\text{suf}}$ and a revised action $\hat{a}^{t}$, thereby enabling online error correction during inference. The correction procedure can be repeated for up to $N_{\text{corr}}$ attempts. Importantly, each attempt only requires regenerating the suffix $\hat{r}^{t}_{\text{suf}}$ and the corresponding action $\hat{a}^{t}$, rather than recomputing the entire trajectory. Thus, the additional computational cost is minimal compared to the total cost.
@@ 3 Experiments
Table 1: Comparison with state-of-the-art methods on mathematical benchmarks, the best results are in bold and the second-best are underlined. Code use indicates whether code tools are employed. †  denotes our reimplementation results of Avg@4, ‡  indicates results from their official releases.